OpenStudy (anonymous):

how do i solve this with ax+by=c Chang the trainer has two solo workout plans that he offers his clients: Plan A and Plan B. Each client does either one or the other (not both). On Friday there were 2 clients who did Plan A and 5 who did Plan B. On Saturday there were 8 clients who did Plan A and 3 who did Plan B. Chang trained his Friday clients for a total of 8 hours and his Saturday clients for a total of 15 hours. How long does each of the workout plans last?

OpenStudy (anonymous):

Chang the trainer has two solo workout plans that he offers his clients: Plan A and Plan B. Each client does either one or the other (not both). On Friday there were 2 clients who did Plan A and 5 who did Plan B. On Saturday there were 8 clients who did Plan A and 3 who did Plan B. Chang trained his Friday clients for a total of 8 hours and his Saturday clients for a total of 15 hours. How long does each of the workout plans last?

563blackghost (563blackghost):

well the equation ax + by = c is standard form.

OpenStudy (anonymous):

@563blackghost would i set it up as 5x+8y=8?

OpenStudy (anonymous):

this ones a lil different from others I've done so I'm kinda stuck on how to set the equation up @563blackghost

OpenStudy (anonymous):

not trying to cheat

OpenStudy (anonymous):

2x+5y=8 3x+8=15 @563blackghost

OpenStudy (anonymous):

*8y srry

OpenStudy (anonymous):

nope not it i got x=-11 and y=6

OpenStudy (anonymous):

@Legends could you help me set this up, i know how to solve

OpenStudy (legends):

im sorry i don't know this one sorry

OpenStudy (anonymous):

np

OpenStudy (anonymous):

@Okean i got x=3/2 y=1